Construct the 99% interval estimate for the ratio of the population variances using the following results from two independently drawn samples from normally distributed populations. (Round P value and final answers to 2 decimal places. You may find it useful to reference the appropriate table: chi-square table or F table) Sample 1: 1 195, si-24.6, and m Sample 2: i2-191.5, 2-221, and n2-8 to <. Prev 40, 5 Next > 344 P
Answer #1

df1=9-1=8

df2=8-1=7

alpha=0.01

F1-α/2 = 8.68 [ excel formula =F.INV.RT(0.005,8,7)]

Fα/2 =0.13 [ excel formula =F.INV(0.005,8,7)]

confidence interval for ration of population variance is
lower bound=   (s1²/s2²)/F1-α/2 =(24.6/22.1)/8.68=0.128267752

upper bound=   (s1²/s2²)/Fα/2 =(24.6/22.1)/0.13 = 8.564521

so, confidence interval is (0.1283 , 8.5645)